A ball of mass 100g strikes a wall with the speed 2m/s and the bounces off the wall in the opposite direction with the same speed calculate the impulse of the ball

1
Expert's answer
2022-02-14T14:23:24-0500

Impulse moment theorem

"\u2206I=\u2206P=m(u-v)"

"u=2m/sec\\v=-2m/sec"

Put value


"∆I=0.1\times(2-(-2))=0.1\times4=0.4kgm /sec"